# Validation of the DESI 2024 Lyman Alpha Forest BAL Masking Strategy

### v1

@@ 2.2 Mocks
The mocks are created in two stages. The first stage is the creation of the transmitted flux skewers for the sight lines to each quasar. This step uses a Gaussian random field to simulate the matter distribution, and the the quasar positions are from the log-normal transformation. The second stage combines those transmission skewers with mock quasar spectra that are representative of the distribution of quasars in DESI. This includes a range of quasar spectral energy distributions and magnitudes, as well as the addition of noise and other astrophysical effects that make the spectra more realistic. These astrophysical effects include metal absorption and BALs. The DESI Ly DR1 analysis uses two types of mocks referred to as LyCoLoRe [55] and Saclay [56] mocks. In this paper we only consider the LyCoLoRe mocks because BALs are added in exactly the same way to both types of mocks. The differences between these two types of mocks include factors such as how the quasar distribution and velocity field are modeled.
@@
The density and velocity distributions for the transmitted flux skewers use the CoLoRe package [57] to generate Gaussian random fields and the Newtonian potential of this field to determine the radial velocity field. We then convert skewers through this density and radial velocity field into skewers of transmitted flux with the LyCoLoRe [55] package. This package adds additional, small scale information based on a one-dimensional Gaussian and computes the optical depth with the fluctuating Gunn-Peterson approximation [58, 59], as well as adds redshift space distortions based on the radial velocity field.
@@
These transmitted flux skewers are added to a quasar population that matches the magnitude, redshift, sky distribution, and density of objects on the sky of the DESI DR1 dataset. One update in DR1 relative to the EDR mocks is a change in the way the mock distribution samples inhomogeneities in the observational data, which is described in detail in [25]. We use the quickquasars111https://github.com/desihub/desisim/blob/main/py/desisim/scripts/quickquasars.py script from the desisim222https://github.com/desihub/desisim package to generate a synthetic quasar spectrum for each quasar, add astrophysical features to make the spectra more realistic, and then add the appropriate level of noise based on the magnitude of each quasar and the number of observations of that quasar available for DR1. The astrophysical features include Damped Lyman systems (DLAs), BALs, and absorption from metals in the IGM, specifically the Si II , Si II , Si III , and Si II lines that are most important absorption features in the forest region. The DLAs and metals are added based on the same density field used to generate the transmitted flux skewers, although the relative strengths of the metal lines are tuned following a procedure described in [25] to match the observational data. The BALs are randomly applied to 16% of quasars.
@@ 3.2 Identification Algorithm
The BAL identification algorithm used for DESI DR1 is nearly identical to the one used by [30] for EDR, so we only briefly summarize it here. The algorithm is based on fitting a series of templates to every spectrum around the rest-frame C IV emission feature and searching for absorption relative to this model fit. This fitting process requires that spectra extend to at least 25,000 km s-1 blueshift relative to the C IV line and at least to 1633 Å in the quasar rest frame in order to fit the quasar spectrum. This sets a redshift limit of based on the observed wavelength range of the DESI data.
@@
We use the best spectral fit to search for BALs associated with C IV based on the AI and BI criteria described in section 3.1. Should a BAL feature be present, we iteratively mask the velocity range of the trough and refit the templates for either ten iterations or convergence. If there is sufficient spectral coverage, we also perform a separate search in the vicinity of the Si IV emission feature. We then record the AI and BI values, velocity range of each trough, and other parameters as described in [30].
@@
We use the public baltools333https://github.com/paulmartini/baltools software package to identify and measure the BALs in DESI. This code was originally developed by [61] to measure the parameters of BALs that were identified via their convolutional neural network, which they applied to SDSS DR14 data. The code was later applied to identify BALs for the DR16 quasar catalog [66], although that worked dropped the CNN component for classification and instead relied on just the measurement of AI and BI to identify BALs. This approach was also adopted by [30] for the DESI EDR BAL catalog, and we continue that approach here.
@@
The most significant change in the baltools package since the EDR catalog is the change from PCA templates calculated by [61] to new quasar templates developed by [67]. These new templates perform better than the old PCA templates, especially for high SNR spectra, and lead to fewer false positives in high SNR data.
@@
We used the baltools package to search for BALs associated with every DESI DR1 quasar in the redshift range . The BAL fraction is 16.7% in the redshift range used for the Ly analysis based on the criterion and 1.3% based on the criterion . These percentages are similar to those measured by [30] based on the first two months of the main survey, although the AI percentage is slightly higher and the BI percentage is slightly lower. The AI and BI distributions of the BALs are shown in figure 2, which also shows divisions of the AI distribution into four quartiles that are separated by , , and . These values are similar although somewhat smaller than the values for the SDSS distribution, which is likely due to the different SNR and perhaps other features of the data.
@@ 4 Impact on Redshift Errors
The impact of BALs on redshift errors was thoroughly studied with mock DESI spectra by [69]. Those authors showed that the redshifts derived by Redrock for mock spectra with BALs had offsets of relative to the same mock quasar spectra without the BAL features. They then masked the locations of pixels that were potentially impacted by BALs, specifically C IV, Si IV, N V, and Ly by setting the inverse variance of pixels to zero if they were at the same velocity relative to each feature as the BAL absorption associated with C IV. The redshift offsets for the masked BALs were substantially smaller than for the unmasked BALs, where in both cases the redshift offsets were relative to the redshift if the BAL features were not present. Those authors found that masking the locations of the BAL features led to a substantial reduction in the redshift errors relative to the true value, as well as improvements in other areas such as the number of catastrophic redshift errors and spectral misclassifications.
@@
We have repeated the masking analysis performed by [30] for the DESI DR1 data to measured the typical velocity differences between masking and not masking BALs, as well as to evaluate the impact of these errors on the BAO parameters. Figure 5 shows the change in velocity for the entire BAL sample, as well as distributions for the four quartiles in AI values. The average and median offsets range from nearly zero for the lowest quartile to km s-1 for the largest quartile. These offsets are substantially smaller than those found by [30], which is likely because the DR1 quasar redshifts were measured with new quasar templates from [52]. That paper refit quasars from the [30] study with their new quasar templates and measured the cross-correlation function between the quasars with BALs and Ly absorbers to measure the average bias in the quasar redshifts. They found that the average shift was km s-1 without masking with the new templates, as compared to km s-1 without masking with the previous quasar templates, which is roughly consistent with the change between the average of km s-1 measured by [30] with EDR data and the previous quasar templates and the average of km s-1 we measure with the DR1 data and the new templates.
@@ 5 Impact on Baryon Acoustic Oscillations
Nearly 17% of the DESI DR1 quasars with are observed to have BAL features associated with C IV based on the AI criterion (see table 1), and these BAL features add absorption in the forest region and impact redshift errors. In this section we evaluate the impact of BALs on the measurement of the location of the BAO peak with a range of alternative masking strategies. There are two motivations for this study. First, the BAL masking strategy we employ is quite conservative, in that we mask pixels in the velocity range of every BAL features identified in the stacking study of [65]. Most of these BAL features are expected to be quite weak, and may be negligible compared to the forest absorption, so the masking strategy may remove pathlength unnecessarily. We consequently fit the BAO peak with a range of alternative strategies that mask fewer BALs. These strategies are to only mask BALs above the AI values that divide each quartile in the data (see figure 2), that is above , , and . In each of these cases we mask BALs with AI values above this threshold and do not mask any features in BALs below this threshold. We also consider the case where we just mask BALs with . The second motivation is that we know the BAL identification algorithm is incomplete based on our studies with mocks, and therefore there are unidentified BALs in our data that may compromise the analysis. We use the mock data to compared the results between catalogs based on the true BAL distribution and the ‘observed’ distribution returned by the identification algorithm.
@@ 5.1 Continuum Fitting
We create alternative catalogs for each masking option by setting the AI and velocity ranges for the BAL features to zero for the quasars that we do not want to mask. For example, for our option where we only mask BALs with , we create a version of the catalog where we set the AI and velocity range values for all BALs with to be equal to zero. For the case where we only mask BALs with , we set AI and the velocity ranges equal to zero for every BAL that has . Note the masking for the BALs that remain are based on the velocity range associated with the criterion, although these largely correspond to the same pixels as the velocity range for their troughs that meet the BI criterion. In addition, we consider one case where we adjust the redshifts for the BALs as described in Section 4. After we adjusted the redshifts, we reran the BAL detection algorithm based on new redshifts. This led to very minor changes in the final catalog of BAL properties, such as an 0.1% change in the BAL fraction, and consequently very minor changes in which pixels were masked. As a reference, we also perform all of our analysis with no BALs masked (“no masking”), although this is just intended as a point of comparison and not a viable alternative strategy. In addition, DLAs are still masked in all of these options. In total, this corresponds to 19 different options: seven different options with DESI data (including the baseline analysis, no masking, and the updated redshifts) and 12 different options on mocks (six with the true BAL catalog, six with the “observed” BAL catalog).
@@
where is the measured flux, is the global mean flux, and is the quasar continuum. Since the forest is too dense at these redshifts to measure the true continuum directly, the package iteratively fits the quasar continuum and from the data:
where is the mean continuum calculated from the full sample, is the rest-frame wavelength, , and are parameters that are fit separately to each quasar to account for spectral diversity. For more details, see [22].
The continua are extremely similar for the different masking options. The ratio of the continuum for each masking option relative to the baseline is only a few tenths of a percent for Ly region A ( Å) and under a percent for Ly region B ( Å). We refer to these two regions as Ly(A) and Ly(B). Unsurprisingly, the biggest change is in comparison to the “no masking” case. Yet even in this case the dispersion in Ly(B) is slightly less than a percent and about half a percent in Ly(A). These differences are much smaller than those shown in [31], as the figures in that paper just showed the continua for the different subsets of BAL quasars.
@@ 5.2 Correlation Function and Uncertainties
We measured all four correlations based on the overdensity fields for each of the nineteen different scenarios described in the previous subsection, that is seven on data and 12 on mocks. The four correlations are the autocorrelation of Ly(A)Ly(A), the autocorrelation of Ly(A)Ly(B), the cross-correlation of Ly(A)QSO, and the cross correlation of Ly(B)QSO. Our analysis of the data closely followed the procedure described in detail in [24] and our analysis of the mocks closely followed [25].

### v2

@@ 2.2 Mocks
The mocks are created in two stages. The first stage is the creation of the transmitted flux skewers for the sight lines to each quasar. This step uses a Gaussian random field to simulate the matter distribution, and the the quasar positions are from the log-normal transformation [e.g. 55]. The second stage combines those transmission skewers with mock quasar spectra that are representative of the distribution of quasars in DESI. This includes a range of quasar spectral energy distributions and magnitudes, as well as the addition of noise and other astrophysical effects that make the spectra more realistic. The noise is added based on a model for instrument that includes the throughput and detector properties and the astrophysical effects include metal absorption and BALs [see 54, 25, for details]. The DESI Ly DR1 analysis uses two types of mocks referred to as LyCoLoRe [56] and Saclay [57] mocks. In this paper we only consider the LyCoLoRe mocks because BALs are added in exactly the same way to both types of mocks. The differences between these two types of mocks include factors such as how the quasar distribution and velocity field are modeled.
@@
The density and velocity distributions for the transmitted flux skewers use the CoLoRe package [58] to generate Gaussian random fields and the Newtonian potential of this field to determine the velocity field for the skewers. We then convert skewers through this density and radial velocity field into skewers of transmitted flux with the LyCoLoRe [56] package. This package adds additional, small scale information based on a one-dimensional Gaussian and computes the optical depth with the fluctuating Gunn-Peterson approximation [59, 60], as well as adds redshift space distortions based on the radial velocity field.
@@
These transmitted flux skewers are added to a quasar population that matches the magnitude, redshift, sky distribution, and density of objects on the sky of the DESI DR1 dataset. One update in DR1 relative to the EDR mocks is a change in the way the mock distribution samples inhomogeneities in the observational data, which is described in detail in [25]. We use the quickquasars111https://github.com/desihub/desisim/blob/main/py/desisim/scripts/quickquasars.py script from the desisim222https://github.com/desihub/desisim package to generate a synthetic quasar spectrum for each quasar, add astrophysical features to make the spectra more realistic, and then add the appropriate level of noise based on the magnitude of each quasar and the number of observations of that quasar available for DR1. The astrophysical features include Damped Lyman systems (DLAs), BALs, and absorption from metals in the IGM, specifically the Si II , Si II , Si III , and Si II lines that are most important absorption features in the forest region. The DLAs and metals are added based on the same density field used to generate the transmitted flux skewers, although the relative strengths of the metal lines are tuned following a procedure described in [25] to match the observational data. The BALs are randomly applied to 16%. This percentage is based on measurements from SDSS and early DESI data [61, 30].
@@ 3.2 Identification Algorithm
The BAL identification algorithm we use for DESI DR1 and mocks is nearly identical to the one used by [30] for EDR, so we only briefly summarize it here. The algorithm is based on fitting a series of templates to every spectrum around the rest-frame C IV emission feature and searching for absorption relative to this model fit. We set the blue limit of this fit to 25,000 km s-1 blueshift (about rest frame 1420 Å) relative to the C IV line. For most quasars the spectral range extends to 2400 Å and we set a minimum of at least to 1633 Å for the highest redshifts. This sets a redshift range of based on the observed wavelength range of the DESI data. We shift the spectrum to the rest frame based on the redshift calculated by Redrock. This redshift uses a prior from one of the other tools described in 2.1 if Redrock did not initially identify the object as a quasar.
@@
We use the best spectral fit to search for BALs associated with C IV based on the AI and BI criteria described in section 3.1, namely with the ratio of the observed flux to the best spectral fit. Should a BAL feature be present, we iteratively mask the velocity range of the trough and refit the templates for either ten iterations or convergence. The iteration process improves the quality of the continuum fit outside of the trough region(s). If there is sufficient spectral coverage, we also perform a separate search in the vicinity of the Si IV emission feature. We then record the AI and BI values, velocity range of each trough, and other parameters as described in [30].
@@
We use the public baltools333https://github.com/paulmartini/baltools software package to identify and measure the BALs in DESI. This code was originally developed by [63] to measure the parameters of BALs that were identified via their convolutional neural network, which they applied to SDSS DR14 data. The code was later applied to identify BALs for the DR16 quasar catalog [68], although that worked dropped the CNN component for classification and instead relied on just the measurement of AI and BI to identify BALs with the algorithm described in this subsection. This approach was also adopted by [30] for the DESI EDR BAL catalog, and we continue that approach here.
@@
The most significant change in the baltools package since the EDR catalog is the change from PCA templates calculated by [63] to new quasar templates developed by [69]. These new templates perform better than the old PCA templates, including the fit to the C 4 region, especially for high SNR spectra, and lead to fewer false positives in high SNR data. For example, based on visual inspection the number of false positives at high SNR (SNR) decreased from about 10% to consistent with zero.
@@
We used the baltools package to search for BALs associated with every DESI DR1 quasar in the redshift range . The BAL fraction is 16.7% in the redshift range used for the Ly analysis based on the criterion and 1.3% based on the criterion . These percentages are similar to those measured by [30] based on the first two months of the main survey, although the AI percentage is slightly higher and the BI percentage is slightly lower. The AI and BI distributions of the BALs are shown in figure 2, which also shows divisions of the AI distribution into four quartiles that are separated by , , and . These values are similar although somewhat smaller than the values for the SDSS distribution [250, 839, and 2221, respectively, see 31], which is likely due to the different SNR and perhaps other features of the data.
@@ 4 Impact on Redshift Errors
The impact of BALs on redshift errors was thoroughly studied with mock DESI spectra by [71]. The redshift error or velocity shift is defined as:
where is the measureded redshift without masking and is the redshift measured after masking the BAL features. [71] showed that the redshifts derived by Redrock for mock spectra with BALs had offsets of relative to the same mock quasar spectra without the BAL features. They then masked the locations of pixels that were potentially impacted by BALs, specifically C IV, Si IV, N V, and Ly by setting the inverse variance of pixels to zero if they were at the same velocity relative to each feature as the BAL absorption associated with C IV. The redshift offsets for the masked BALs were substantially smaller than for the unmasked BALs, where in both cases the redshift offsets were relative to the redshift if the BAL features were not present. Those authors found that masking the locations of the BAL features led to a substantial reduction in the redshift errors relative to the true value, as well as improvements in other areas such as the number of catastrophic redshift errors and spectral misclassifications.
@@
We have repeated the masking analysis performed by [30] for the DESI DR1 data to measure the typical velocity differences between masking and not masking BALs, as well as to evaluate the impact of these errors on the BAO parameters. Figure 5 shows the change in velocity for the entire BAL sample, as well as distributions for the four quartiles in AI values. The average and median offsets range from nearly zero for the lowest quartile to km s-1 for the largest quartile. These offsets are substantially smaller than those found by [30], which is likely because the DR1 quasar redshifts were measured with new quasar templates from [52]. That paper refit quasars from the [30] study with their new quasar templates and measured the cross-correlation function between the quasars with BALs and Ly absorbers to measure the average bias in the quasar redshifts. They found that the average shift was km s-1 without masking with the new templates, as compared to km s-1 without masking with the previous quasar templates, which is roughly consistent with the change between the average of km s-1 measured by [30] with EDR data and the previous quasar templates and the average of km s-1 we measure with the DR1 data and the new templates.
@@ 5 Impact on Baryon Acoustic Oscillations
Nearly 17% of the DESI DR1 quasars with are observed to have BAL features associated with C IV based on the AI criterion (see table 1), and these BAL features add absorption in the forest region and impact redshift errors. In this section we evaluate the impact of BALs on the measurement of the location of the BAO peak with a range of alternative masking strategies. There are two motivations for this study. First, the BAL masking strategy we employ is quite conservative, in that we mask pixels in the velocity range of every BAL features identified in the stacking study of [67]. Most of these BAL features are expected to be quite weak, and may be negligible compared to the forest absorption, so the masking strategy may remove pathlength unnecessarily. We consequently fit the BAO peak with a range of alternative strategies that mask fewer BALs. These strategies are to only mask BALs above the AI values that divide each quartile in the data (see figure 2), that is above , , and . In each of these cases we mask BALs with AI values above this threshold and do not mask any features in BALs below this threshold. We also consider the case where we just mask BALs with , as well as show the case (baseline) and (masked z), which has updated redshifts as described in Section 4. The second motivation is that we know the BAL identification algorithm is incomplete based on our studies with mocks, and therefore there are unidentified BALs in our data that may compromise the analysis. We use the mock data to compared the results between catalogs based on the true BAL distribution and the ‘observed’ distribution returned by the identification algorithm.
@@ 5.1 Continuum Fitting
We create alternative catalogs for each masking option by setting the AI and velocity ranges for the BAL features to zero for the quasars that we do not want to mask. For example, for our option where we only mask BALs with , we create a version of the catalog where we set the AI and velocity range values for all BALs with to be equal to zero. For the case where we only mask BALs with , we set AI and the velocity ranges equal to zero for every BAL that has . Note the masking for the BALs that remain are based on the velocity range associated with the criterion, although these largely correspond to the same pixels as the velocity range for their troughs that meet the BI criterion. In addition, we consider one case where we adjust the redshifts for the BALs as described in Section 4. After we adjusted the redshifts, we reran the BAL detection algorithm based on new redshifts. This led to very minor changes in the final catalog of BAL properties, such as an 0.1% change in the BAL fraction, and consequently very minor changes in which pixels were masked. We refer to this case as “ (masked z).” As a reference, we also perform all of our analysis with no BALs masked (“no masking”), although this is just intended as a point of comparison and not a viable alternative strategy. In addition, DLAs are still masked in all of these options. In total, this corresponds to 19 different options: seven different options with DESI data (including the baseline analysis, no masking, and the updated redshifts) and 12 different options on mocks (six with the true BAL catalog, six with the “observed” BAL catalog).
@@
where is the measured flux, is the mean transmission of the intergalactic medium, and is the mean quasar continuum. Since the forest is too dense at these redshifts to measure the true continuum directly, the package iteratively fits the quasar continuum and from the data:
where is the mean quasar continuum calculated from the full sample, is the rest-frame wavelength, , and are parameters that are fit separately to each quasar to account for spectral diversity. For more details, see [22].
The continua are extremely similar for the different masking options described at the beginning of this section. The ratio of the continuum for each masking option relative to the baseline is only a few tenths of a percent for Ly region A ( Å) and under a percent for Ly region B ( Å). We refer to these two regions as Ly(A) and Ly(B). Unsurprisingly, the biggest change is in comparison to the “no masking” case. Yet even in this case the dispersion in Ly(B) is slightly less than a percent and about half a percent in Ly(A). These differences are much smaller than those shown in figures 2 – 4 of [31], as the figures in that paper just showed the continua for the different subsets of BAL quasars.
@@ 5.2 Correlation Function and Uncertainties
We measured all four correlations based on the overdensity fields for each of the nineteen different scenarios described in the previous subsection, that is the seven on data listed in Table 3 and 12 on mocks (six with the truth catalog, six with the catalog from the detection algorithm). The four correlations are the autocorrelation of Ly(A)Ly(A), the autocorrelation of Ly(A)Ly(B), the cross-correlation of Ly(A)QSO, and the cross correlation of Ly(B)QSO. Our analysis of the data closely followed the procedure described in detail in [24] and our analysis of the mocks closely followed [25].